Circular No. 6675

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE NOVA W. Liller, Viña del Mar, Chile, reports his discovery of an apparent

nova (mag 8.5) on a Technical Pan photographs taken with Problicom on June 5.085 and 5.087 UT with an 85-mm camera lens (+ orange filter), located at α = 17^{h}54^{m}.2, δ = −30^{o}02′ (equinox 2000.0). Nothing brighter than mag 11.5 appears at this location on a photograph taken on June 2.09. CCD magnitude on June 6.120: V = 9.0 (GSC comparison stars).

SUPERNOVA 1997cq IN ANONYMOUS GALAXY S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrogram of SN 1997cq (cf. IAUC 6674), taken by P. Berlind on June 5.34 UT at the Whipple Observatory 1.5-m telescope, shows it to be a type-IIn supernova at a redshift of z = 0.032. The spectrum shows narrow Hα emission, characteristic of type IIn, with a broader underlying component of width 2800 km/s FWHM. Images taken concurrently with the 1.2-m telescope by L. Macri and D. Koranyi yield magnitudes V = 19.0 and R = 18.9.

A. V. Filippenko and D. C. Leonard, University of California at Berkeley, report that CCD spectra (range 320–1000 nm) obtained with the Lick 3-m Shane reflector on June 6 show that SN 1997cq is a peculiar type-II supernova (formally known as type IIn), dominated by relatively narrow (FWHM < 1000 km/s) Balmer emission lines on a featureless continuum.

COMET C/1996 J1 (EVANS-DRINKWATER) Extension to ephemeris on IAUC 6653 (elements MPC 29879):

1997 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 11 2^{h}05.^{m}39 +31^{o}04.^{′}4 3.249 2.602 43.^{o}1 15.^{o}5 13.2
21 2 19.82 +32 38.6 3.270 2.709 48.6 16.4 13.4
July 1 2 32.89 +34 04.8 3.280 2.815 54.6 17.1 13.6
11 2 44.47 +35 23.9 3.277 2.922 60.9 17.7 13.7
21 2 54.39 +36 36.4 3.264 3.028 67.7 18.1 13.9
31 3 02.49 +37 42.6 3.242 3.133 74.8 18.2 14.0
Aug. 10 3 08.56 +38 42.5 3.212 3.238 82.4 18.1 14.1
20 3 12.41 +39 35.5 3.178 3.343 90.4 17.6 14.3
1997 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green